OpenStudy (amberlykhan):

Solve using the quadratic formula. 1. x^2 + x + 12 2. 4x^2 - 17x - 15 = 0

OpenStudy (astrophysics):

Use \[\huge x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (astrophysics):

Where your expression is of the sort \[ax^2+by+c =0\]

OpenStudy (amberlykhan):

Im stumped. So far I got x = -1 +/- (square of -47) all divided by 2

OpenStudy (amberlykhan):

@Astrophysics

OpenStudy (amberlykhan):

what do i do now

OpenStudy (astrophysics):

\[x^2+x+12 \implies a=1,b =1, c = 12\] \[x = \frac{ -1\pm \sqrt{1^2-4(1)(12)} }{ 2(1) } = \frac{ -1 \pm \sqrt{-47} }{ 2 }\] this requires complex numbers, do you know them?

OpenStudy (anonymous):

stick to your guns it is right unless you want to use \(i\)

OpenStudy (astrophysics):

\[x = \frac{ -1\pm \sqrt{1^2-4(1)(12)} }{ 2(1) } = \frac{ -1 \pm \sqrt{-47} }{ 2 } = \frac{ -1 \pm i \sqrt{47} }{ 2 }\]

OpenStudy (astrophysics):

This lag is insane

OpenStudy (astrophysics):

Hey, yeah do as satellite says if you don't know complex numbers leave it, I'm sure your teacher doesn't want you to use it, but in any case \[i = \sqrt{-1}\]

OpenStudy (sbuck98):

Do you need anymore help?